# Flips in colorful triangulations

### v1

@@ Abstract.
The associahedron is the graph ${\mathcal{G}}_{N}$ that has as nodes all triangulations of a convex $N$-gon, and an edge between any two triangulations that differ in a flip operation, which consists of removing an edge shared by two triangles and replacing it by the other diagonal of the resulting 4-gon. In this paper, we consider a large collection of induced subgraphs of ${\mathcal{G}}_{N}$ obtained by Ramsey-type colorability properties. Specifically, coloring the points of the $N$-gon red and blue alternatingly, we consider only colorful triangulations, namely triangulations in which every triangle has points in both colors, i.e., monochromatic triangles are forbidden. The resulting induced subgraph of ${\mathcal{G}}_{N}$ on colorful triangulations is denoted by ${\mathcal{F}}_{N}$. We prove that ${\mathcal{F}}_{N}$ has a Hamilton cycle for all $N\geq 8$, resolving a problem raised by Sagan, i.e., all colorful triangulations on $N$ points can be listed so that any two cyclically consecutive triangulations differ in a flip. In fact, we prove that for an arbitrary fixed coloring pattern of the $N$ points with at least 10 changes of color, the resulting subgraph of ${\mathcal{G}}_{N}$ on colorful triangulations (for that coloring pattern) admits a Hamilton cycle. We also provide an efficient algorithm for computing a Hamilton path in ${\mathcal{F}}_{N}$ that runs in time ${\mathcal{O}}(1)$ on average per generated node. This algorithm is based on a new and algorithmic construction of a tree rotation Gray code for listing all $n$-vertex $k$-ary trees that runs in time ${\mathcal{O}}(k)$ on average per generated tree.
@@ 1. Introduction
The associahedron is a polytope of fundamental interest and importance [Lod04, HL07, CSZ15]. In this paper we are specifically interested in its combinatorial structure, namely the graph of its skeleton; see Figure 1. This graph, which we denote by ${\mathcal{G}}_{N}$, has as nodes all triangulations of a convex $N$-gon ($N\geq 3$), and an edge between any two triangulations that differ in a flip operation, which consists of removing an edge shared by two triangles and replacing it by the other diagonal of the resulting 4-gon. The graph ${\mathcal{G}}_{N}$ is isomorphic to the graph that has as nodes all binary trees with $N-2$ vertices, and an edge between any two trees that differ in a tree rotation. Each binary trees arises as the geometric dual of a triangulation, with the root given by ‘looking through’ a fixed outer edge, and flips translate to tree rotations under this bijection; see Figure 2. This dual point of view between triangulations and binary trees, and more generally, between dissections into $(k+1)$-gons and $k$-ary trees, is an essential tool used in our paper.
@@
Another fundamental graph property that we focus on in this paper is Hamiltonicity. To this end, Lucas [Luc87] first proved that ${\mathcal{G}}_{N}$ admits a Hamilton cycle for $N\geq 5$, and a short proof was given by Hurtado and Noy [HN99]. A Hamilton path in ${\mathcal{G}}_{N}$ can be computed efficiently and yields a Gray code ordering of all binary trees by rotations [LRvBR93]. This algorithm is a special case of the more general Hartung-Hoang-Mütze-Williams permutation language framework [HHMW22, HM21, MM23, CMM22].
In this paper, we consider a large collection of induced subgraphs of ${\mathcal{G}}_{N}$ obtained by Ramsey-type colorability properties, a line of inquiry initiated by Propp and Sagan [Sag08]. Specifically, we label the points of the convex $N$-gon by $1,\ldots,N$ in counterclockwise order, and we color them red (r) and blue (b) alternatingly. It follows that point $i$ is colored red if $i$ is odd and blue if $i$ is even. Note that for even $N$, any two neighboring points have opposite colors, whereas for odd $N$ this property is violated for the first and last point, which are both red.
@@ 1.3. Three colors
Figure 8. Twist operation and the corresponding binary trees.
Figure 9. The flip graphs ${\mathcal{H}}_{6}$ and ${\mathcal{H}}_{9}$.
@@ 2.2. Dissections and trees
For any coloring sequence $\alpha$, we write ${\mathcal{C}}_{\alpha}$ for the set of colorful triangulations with coloring pattern (1). By these definitions, ${\mathcal{F}}_{\alpha}$ is the subgraph of ${\mathcal{G}}_{N}$ induced by the triangulations in ${\mathcal{C}}_{\alpha}$. Sagan’s question concerned the special case $\alpha:=1^{N}$ for even $N$ and $\alpha:=21^{N-2}$ for odd $N$, and for those particular coloring sequences $\alpha$ we simply write ${\mathcal{C}}_{N}={\mathcal{C}}_{\alpha}$ and ${\mathcal{F}}_{N}={\mathcal{F}}_{\alpha}$. Sagan proved the following.
@@ 2.4. Graphs
We refer to the vertices $0^{a}$ and $d^{a}$ as extremal vertices, and note that they have degree 1 in $G(a,d)$, unless $a=d=1$, in which case the graph is a single vertex having degree 0.
@@ 3. Colorful triangulations
In this section we turn to the setting of colorful triangulations introduced by Sagan, with the goal of proving Theorems 1 and 2.
@@ 3.1. Alternating colors
    Every boundary edge of $T$ is colorful.
@@
    Every monochromatic edge of $T$ is an inner edge, and the other four edges bounding the two triangles adjacent to this edge are all colorful.
  3. (iii)
    Every triangle of $T$ has one monochromatic edge.
Consequently, if we remove from $T$ all monochromatic edges, keeping only the colorful ones, then the resulting dissection $r(T)$ is a quadrangulation on the point set. Indeed, by (i)+(iii) every triangle is destroyed, and by (ii) destroying a triangle creates a quadrangle. While $T$ has $N-2=n$ triangles, $r(T)$ has $q:=(N-2)/2=n/2$ quadrangles. Furthermore, there are $2^{q}$ many colorful triangulations that yield the same quadrangulation $r(T)$ by removing monochromatic edges. They are obtained from $r(T)$ by placing a diagonal in each of the $q$ quadrangles in one of the two ways. Note that the subgraph of ${\mathcal{F}}_{N}$ induced by those $2^{q}$ triangulations is isomorphic to the $q$-dimensional hypercube $Q_{q}$, as each of the $q$ monochromatic edges in $T$ can be flipped independently from the others. We thus obtain a partition of ${\mathcal{F}}_{N}$ into hypercubes $Q_{q}$, plus edges between them. These copies of hypercubes are highlighted by blue bubbles in Figure 12.
We also note that every quadrangulation $R$ on $N$ points equals $r(T)$ for some colorful triangulation $T\in{\mathcal{C}}_{N}$. Indeed, given $R$, then coloring the $N$ points red and blue alternatingly will result in the endpoints of every edge receiving both colors, i.e., all edges are colorful. We define a reduced graph ${\mathcal{F}}_{N}^{\prime}$, that has as nodes all quadrangulations on $N$ points, and for any two colorful triangulations $T$ and $T^{\prime}$ that differ in a flip of a colorful edge, we add an edge between $r(T)$ and $r(T^{\prime})$ in ${\mathcal{F}}_{N}^{\prime}$; see Figures 12 and 13. We observe that ${\mathcal{F}}_{N}^{\prime}$ is isomorphic to the flip graph of quadrangulations ${\mathcal{G}}_{N,4}$, i.e., we have ${\mathcal{F}}_{N}^{\prime}\simeq{\mathcal{G}}_{N,4}$. These arguments yield a direct combinatorial proof for the first equality in Theorem 10.
@@
The statement of the lemma is false for $q=2$: Indeed, while ${\mathcal{F}}_{6}^{\prime}$ has a Hamilton path (the graph is a triangle), there is no Hamilton path or cycle in ${\mathcal{F}}_{6}$; see Figure 12.
@@
We now describe both steps in detail. For a given quadrangulation $R$, we associate each of the colorful triangulations $T\in C(R)$ by a bitstring $b(T)\in\{0,1\}^{q}$ as follows: We label the $q$ quadrangles of $R$ arbitrarily by $j=1,\ldots,q$, and we define $b(T)_{j}=0$ or $b(T)_{j}=1$ if the monochromatic edge of $T$ that sits inside the $j$th quadrangle of $R$ connects the two red points or the two blue points, respectively.
@@
If the node $R$ has degree $d$ in ${\mathcal{S}}$, then we have to choose $d$ distinct shortcut edges in the hypercube $Q(R)$, each selected from a distinct (but not necessarily disjoint) $(q-2)$-dimensional subcube, and to find a Hamilton cycle in $Q(R)$ that contains all of these edges; see Figure 14 (b). By Lemma 11, it is enough to ensure that the shortcut edges together form paths in $Q(R)$. If $d\leq 2$ (case (i) of the lemma), then this is clear, as one or two edges always form one or two paths. If $d=3$ (case (ii) of the lemma), one has to avoid that all three shortcut edges are incident to the same node, which is easily possible under the stronger assumption $q\geq 4$.
Then the Hamilton cycle in ${\mathcal{F}}_{N}$ is obtained by taking the symmetric difference of the edge sets of the Hamilton cycles in $Q(R)$ for all quadrangulations $R$ on $N$ points with the 4-cycles formed by the connectors and shortcuts (i.e., the shortcuts are removed, and the connectors are added instead). This completes the proof. ∎
@@ 3.2. General coloring patterns
We now consider an arbitrary coloring sequence $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ and the corresponding coloring pattern defined in (1). Recall that ${\mathcal{C}}_{\alpha}$ denotes the set of all colorful triangulations with this coloring pattern. The corresponding flip graph is denoted by ${\mathcal{F}}_{\alpha}$, and it is an induced subgraph of the associahedron ${\mathcal{G}}_{N}$, where $N=\sum_{i=1}^{\ell}\alpha_{i}$. As in the previous section, a colorful triangulation $T\in{\mathcal{C}}_{\alpha}$ has two types of edges, namely monochromatic and colorful edges. We write $E_{\alpha}$ for the set of boundary edges that are monochromatic in $T$, i.e., these are the pairs of points $(i,i+1)$ for $i=1,\ldots,N$ (modulo $N$) where both endpoints receive the same color. Generalizing the discussion from the previous section, we observe the following:
@@
    Except the edges in $E_{\alpha}$, every boundary edge of $T$ is colorful.
@@
    Except the edges in $E_{\alpha}$, every monochromatic edge of $T$ is an inner edge, and the other four edges bounding the two triangles adjacent to this edge are all colorful.
  3. (iii)
    Every triangle of $T$ has one monochromatic edge.
@@
We refer to a dissection of a convex $N$-gon into $q$ quadrangles and $t$ triangles that contain all the edges of $E_{\alpha}$ as an $\alpha$-angulation, and we write ${\mathcal{D}}_{\alpha}$ for the set of all such dissections. We also note that every $\alpha$-angulation $R$ on $N$ points equals $r(T)$ for some colorful triangulation $T\in{\mathcal{C}}_{\alpha}$. Indeed, given $R$, then coloring the $N$ points according to the pattern (1), the endpoints of every edge except the ones in $E_{\alpha}$ receive both colors, i.e., all edges except the ones in $E_{\alpha}$ are colorful. We define a reduced graph ${\mathcal{F}}_{\alpha}^{\prime}$ that has as nodes all $\alpha$-angulations on $N$ points, and for any two colorful triangulations $T$ and $T^{\prime}$ that differ in a flip of a colorful edge, we add an edge between $r(T)$ and $r(T^{\prime})$ in ${\mathcal{F}}_{\alpha}^{\prime}$; see Figures 12, 13 and 15.
@@
The proof of Lemma 13 presented in the previous section generalizes straightforwardly, yielding the following statement. Note that the assumptions $\ell=8$ and $\ell\geq 10$ in the following lemma translate to $q=(\ell-2)/2=3$ and $q\geq 4$ used in the proof of Lemma 13, respectively.
@@ 4. Proof of Theorem 5
For $a=1$ it is easy to see that there is no Hamilton path. For $a=2$ a Hamilton path is shown in Figure 12 (in the graph ${\mathcal{F}}_{7}$). For $a=3$ a computer search shows that there is no Hamilton path. For $a\geq 4$, a Hamilton path can be constructed by combining the gadgets shown in Figure 19.
@@ 5.1. Tree rotation
We use the following compact notation for $k$-ary trees, defined inductively as follows: The unique empty tree is denoted by $\varepsilon$ (it has 0 vertices). If $T_{1},\ldots,T_{k}$ are $k$-ary trees, then $T:=[T_{1},\ldots,T_{k}]$ is a $k$-ary tree. Furthermore, the subtrees of $T$ are given by $T$ itself and by the subtrees of $T_{i}$ for $i=1,\ldots,k$. If $T_{i}=\cdots=T_{j}=\varepsilon$, then we write $T=[T_{1},\ldots,T_{k}]$ as $[T_{1},\ldots,T_{i-1},\varepsilon^{j-i+1},T_{j+1},\ldots,T_{k}]$. In particular, $[\varepsilon^{k}]$ is the unique 1-vertex $k$-ary tree.

### v2

@@ Abstract.
The associahedron is the graph ${\mathcal{G}}_{N}$ that has as nodes all triangulations of a convex $N$-gon, and an edge between any two triangulations that differ in a flip operation. A flip removes an edge shared by two triangles and replaces it by the other diagonal of the resulting 4-gon. In this paper, we consider a large collection of induced subgraphs of ${\mathcal{G}}_{N}$ obtained by Ramsey-type colorability properties. Specifically, coloring the points of the $N$-gon red and blue alternatingly, we consider only colorful triangulations, namely triangulations in which every triangle has points in both colors, i.e., monochromatic triangles are forbidden. The resulting induced subgraph of ${\mathcal{G}}_{N}$ on colorful triangulations is denoted by ${\mathcal{F}}_{N}$. We prove that ${\mathcal{F}}_{N}$ has a Hamilton cycle for all $N\geq 8$, resolving a problem raised by Sagan, i.e., all colorful triangulations on $N$ points can be listed so that any two cyclically consecutive triangulations differ in a flip. In fact, we prove that for an arbitrary fixed coloring pattern of the $N$ points with at least 10 changes of color, the resulting subgraph of ${\mathcal{G}}_{N}$ on colorful triangulations (for that coloring pattern) admits a Hamilton cycle. We also provide an efficient algorithm for computing a Hamilton path in ${\mathcal{F}}_{N}$ that runs in time ${\mathcal{O}}(1)$ on average per generated node. This algorithm is based on a new and algorithmic construction of a tree rotation Gray code for listing all $n$-vertex $k$-ary trees that runs in time ${\mathcal{O}}(k)$ on average per generated tree.
An extended abstract of this work was accepted for presentation at the 32nd International Symposium on Graph Drawing and Network Visualization (Graph Drawing 2024).
@@ 1. Introduction
The associahedron is a polytope of fundamental interest and importance [Lod04, HL07, CSZ15], as it lies at the heart of many recent developments in algebraic combinatorics and discrete geometry; see [PSZ23] and the references therein. In this paper we are specifically interested in its combinatorial structure, namely the graph of its skeleton; see Figure 1. This graph, which we denote by ${\mathcal{G}}_{N}$, has as nodes all triangulations of a convex $N$-gon ($N\geq 3$), and an edge between any two triangulations that differ in a flip operation, which consists of removing an edge shared by two triangles and replacing it by the other diagonal of the resulting 4-gon. The graph ${\mathcal{G}}_{N}$ is isomorphic to the graph that has as nodes all binary trees with $N-2$ vertices, and an edge between any two trees that differ in a tree rotation. Each binary tree arises as the geometric dual of a triangulation, with the root given by ‘looking through’ a fixed outer edge, and flips translate to tree rotations under this bijection; see Figure 2. This dual point of view between triangulations and binary trees, and more generally, between dissections into $(k+1)$-gons and $k$-ary trees, is an essential tool used in our paper.
@@
Another fundamental graph property that we focus on in this paper is Hamiltonicity. To this end, Lucas [Luc87] first proved that ${\mathcal{G}}_{N}$ admits a Hamilton cycle for $N\geq 5$, and a short proof was given by Hurtado and Noy [HN99]. A Hamilton path in ${\mathcal{G}}_{N}$ can be computed efficiently and yields a Gray code ordering of all binary trees by rotations [LRvBR93]. This algorithm is a special case of the more general Hartung-Hoang-Mütze-Williams permutation language framework [HHMW22, HM21, MM23, CMM24].
In this paper, we consider a large collection of induced subgraphs of ${\mathcal{G}}_{N}$ obtained by Ramsey-type colorability properties. This line of inquiry was initiated by Sagan [Sag08], following a sequence of problems posed by Propp on a mailing list in 2003. Specifically, we label the points of the convex $N$-gon by $1,\ldots,N$ in counterclockwise order, and we color them red (r) and blue (b) alternatingly. It follows that point $i$ is colored red if $i$ is odd and blue if $i$ is even. Note that for even $N$, any two neighboring points have opposite colors, whereas for odd $N$ this property is violated for the first and last point, which are both red.
@@ 1.3. Three colors
Figure 8. Twist operation and the corresponding binary trees.
Figure 9. The flip graphs ${\mathcal{H}}_{6}$ and ${\mathcal{H}}_{9}$.
@@ 2.2. Dissections and trees
For any coloring sequence $\alpha$, we write ${\mathcal{C}}_{\alpha}$ for the set of colorful triangulations with coloring pattern defined in (1). By these definitions, ${\mathcal{F}}_{\alpha}$ is the subgraph of ${\mathcal{G}}_{N}$ induced by the triangulations in ${\mathcal{C}}_{\alpha}$. Sagan’s question concerned the special case $\alpha:=1^{N}$ for even $N$ and $\alpha:=(2,1^{N-2})$ for odd $N$, and for those particular coloring sequences $\alpha$ we simply write ${\mathcal{C}}_{N}={\mathcal{C}}_{\alpha}$ and ${\mathcal{F}}_{N}={\mathcal{F}}_{\alpha}$. Sagan proved the following.
@@ 2.4. Graphs
We refer to the vertices $1^{a}$ and $d^{a}$ as extremal vertices, and note that they have degree 1 in $G(a,d)$, unless $a=d=1$, in which case the graph is a single vertex having degree 0.
@@ 3. Colorful triangulations
In this section we consider the setting of colorful triangulations introduced by Sagan, with the goal of proving Theorems 1 and 2.
@@ 3.1. Alternating colors
    Every triangle of $T$ has exactly one monochromatic edge.
@@
    Every monochromatic edge of $T$ is an inner edge.
Consequently, if we remove from $T$ all monochromatic edges, keeping only the colorful ones, then the resulting dissection $r(T)$ is a quadrangulation on the point set. Indeed, by (i) every triangle is destroyed, and by (i)+(ii) destroying a triangle creates a quadrangle. While $T$ has $N-2=n$ triangles, $r(T)$ has $q:=(N-2)/2=n/2$ quadrangles. Furthermore, there are $2^{q}$ many colorful triangulations that yield the same quadrangulation $r(T)$ by removing monochromatic edges. They are obtained from $r(T)$ by placing a diagonal in each of the $q$ quadrangles in one of the two ways. Note that the subgraph of ${\mathcal{F}}_{N}$ induced by those $2^{q}$ triangulations is isomorphic to the $q$-dimensional hypercube $Q_{q}$, as each of the $q$ monochromatic edges in $T$ can be flipped independently from the others. We thus obtain a partition of ${\mathcal{F}}_{N}$ into hypercubes $Q_{q}$, plus edges between them. These copies of hypercubes are highlighted by blue bubbles in Figure 12.
We also note that every quadrangulation $R$ on $N$ points equals $r(T)$ for some colorful triangulation $T\in{\mathcal{C}}_{N}$. Indeed, given $R$, then coloring the $N$ points red and blue alternatingly will make all edges colorful. We define a reduced graph ${\mathcal{F}}_{N}^{\prime}$, that has as nodes all quadrangulations on $N$ points, and for any two colorful triangulations $T$ and $T^{\prime}$ that differ in a flip of a colorful edge, we add an edge between $r(T)$ and $r(T^{\prime})$ in ${\mathcal{F}}_{N}^{\prime}$; see Figures 12 and 13. We observe that ${\mathcal{F}}_{N}^{\prime}$ is isomorphic to the flip graph of quadrangulations ${\mathcal{G}}_{N,4}$, i.e., we have ${\mathcal{F}}_{N}^{\prime}\simeq{\mathcal{G}}_{N,4}$. These arguments yield a direct combinatorial proof for the first equality in Theorem 10.
@@
The conclusions of the lemma do not hold when $q=2$: Indeed, while ${\mathcal{F}}_{6}^{\prime}$ has a Hamilton path (the graph is a triangle), there is no Hamilton path or cycle in ${\mathcal{F}}_{6}$; see Figure 12.
@@
We now describe both steps in detail. For a given quadrangulation $R$, we associate each of the colorful triangulations $T\in C(R)$ by a bitstring $b(T)\in\{0,1\}^{q}$ as follows: We label the $q$ quadrangles of $R$ arbitrarily by $j=1,\ldots,q$, and we define $b(T)_{j}:=0$ if the monochromatic edge of $T$ that sits inside the $j$th quadrangle of $R$ connects the two red points, and otherwise (if it connects two blue points) $b(T)_{j}:=1$.
@@
If the node $R$ has degree $d$ in ${\mathcal{S}}$, then we have to choose $d$ distinct shortcut edges in the hypercube $Q(R)$, each selected from a distinct (but not necessarily disjoint) $(q-2)$-dimensional subcube, and to find a Hamilton cycle $H(R)$ in $Q(R)$ that contains all of these edges; see Figure 14 (b). By Lemma 11, it is enough to ensure that the shortcut edges together form paths in $Q(R)$. If $d\leq 2$ (case (i) of the lemma), then this is clear, as one or two edges always form one or two paths. If $d=3$ (case (ii) of the lemma), one has to avoid that all three shortcut edges are incident to the same node, which is easily possible under the stronger assumption $q\geq 4$.
Then the Hamilton cycle in ${\mathcal{F}}_{N}$ is obtained by taking the symmetric difference of the edge sets of the cycles $H(R)\subseteq Q(R)$ for all quadrangulations $R$ on $N$ points with the 4-cycles formed by the connectors and shortcuts (i.e., the shortcuts are removed, and the connectors are added instead). This completes the proof. ∎
@@ 3.2. General coloring patterns
We now consider an arbitrary coloring sequence $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ and the corresponding coloring pattern defined in (1). Recall that ${\mathcal{C}}_{\alpha}$ denotes the set of all colorful triangulations with this coloring pattern, and that the corresponding flip graph is denoted by ${\mathcal{F}}_{\alpha}$. The graph ${\mathcal{F}}_{\alpha}$ is an induced subgraph of the associahedron ${\mathcal{G}}_{N}$, where $N=\sum_{i=1}^{\ell}\alpha_{i}$. As in the previous section, a colorful triangulation $T\in{\mathcal{C}}_{\alpha}$ has two types of edges, namely monochromatic and colorful edges. We write $E_{\alpha}$ for the set of boundary edges that are monochromatic in $T$, i.e., these are the pairs of points $(i,i+1)$ for $i=1,\ldots,N$ (modulo $N$) where both endpoints receive the same color. Generalizing the discussion from the previous section, we observe the following:
@@
    Every triangle of $T$ has exactly one monochromatic edge.
@@
    Except the edges in $E_{\alpha}$, every monochromatic edge of $T$ is an inner edge.
@@
We refer to a dissection of a convex $N$-gon into $q$ quadrangles and $t$ triangles that contain all the edges of $E_{\alpha}$ as an $\alpha$-angulation, and we write ${\mathcal{D}}_{\alpha}$ for the set of all such dissections. We also note that every $\alpha$-angulation $R$ on $N$ points equals $r(T)$ for some colorful triangulation $T\in{\mathcal{C}}_{\alpha}$. Indeed, given $R$, then coloring the $N$ points according to the pattern in (1) will make all edges except the ones in $E_{\alpha}$ colorful. We define a reduced graph ${\mathcal{F}}_{\alpha}^{\prime}$ that has as nodes all $\alpha$-angulations on $N$ points, and for any two colorful triangulations $T$ and $T^{\prime}$ that differ in a flip of a colorful edge, we add an edge between $r(T)$ and $r(T^{\prime})$ in ${\mathcal{F}}_{\alpha}^{\prime}$; see Figures 12, 13 and 15.
@@
The proof of Lemma 13 presented in the previous section generalizes straightforwardly, yielding the following statement. Note that the variable $\ell$ in Lemma 14 below plays the role of $N=2q+2$ in Lemma 13, and so the assumptions $\ell=8$ and $\ell\geq 10$ translate to $q=(\ell-2)/2=3$ and $q\geq 4$ used in the proof of Lemma 13, respectively.
@@ 4. Proof of Theorem 5
For $a=1$ it is easy to see that there is no Hamilton path; see the top left part of Figure 18 or Figure 12 (the graph ${\mathcal{F}}_{6}$). For $a=2$ a Hamilton path is shown in Figure 12 (in the graph ${\mathcal{F}}_{7}$). For $a=3$ a computer search shows that there is no Hamilton path. For $a\geq 4$, a Hamilton path can be constructed by combining the gadgets shown in Figure 19.
@@ 5.1. Tree rotation
We use the following compact notation for $k$-ary trees, defined inductively as follows: The unique empty tree is denoted by $\varepsilon$ (it has 0 vertices). If $T_{1},\ldots,T_{k}$ are $k$-ary trees, then $T:=[T_{1},\ldots,T_{k}]$ represents the $k$-ary tree that has a root vertex with $T_{1},\ldots,T_{k}$ as its subtrees. Furthermore, the subtrees of $T$ are given by $T$ itself and by the subtrees of $T_{i}$ for $i=1,\ldots,k$. If $T_{i}=\cdots=T_{j}=\varepsilon$, then we write $T=[T_{1},\ldots,T_{k}]$ as $[T_{1},\ldots,T_{i-1},\varepsilon^{j-i+1},T_{j+1},\ldots,T_{k}]$. In particular, $[\varepsilon^{k}]$ is the unique 1-vertex $k$-ary tree.